# Non-coherent evolution of closed weakly interacting system leads to equidistribution of probabilities of microstates

### v5

@@ I Introduction
Another notable approach is the study of quantum entanglement and its relationship to thermodynamic irreversibility. Some researchers have proposed that the generation and disruption of entanglement within a quantum system may play a key role in the emergence of macroscopic irreversibility [27, 28].
In the textbook [14], the second law of thermodynamics is derived from an equation known as the Fermi master equation. However, the equation itself is noted to lack a proper derivation. A modern justification for this idea is provided by the Lindblad master equation [15], which describes systems weakly coupled to their environments. In particular, the Fermi master equation is regarded as a Born-Markov approximation of the Lindblad equation [16]. These equations have been shown to exhibit irreversible behavior, even when the underlying microscopic dynamics are unitary. This approach has been applied to various models in quantum thermodynamics [17, 18]. A particularly important result of this line of research is the identification of conditions under which systems reach a steady state resembling the classical Boltzmann distribution [19, 20].
Analogously, in Ref. [29], the authors derive the second law of thermodynamics from a quantum-mechanical analysis of a total system partitioned into an object and its environment. This type of theory has attracted considerable attention in recent years [30, 31, 32]. Additionally, the theory of non-Markovian quantum evolution [33, 34, 35] has garnered significant interest. It focuses on cases where the system’s dynamics are influenced by interactions with the external world.
@@
After performing the averaging over arguments, we obtain the equation of non-coherent evolution. We find that it coincides with the previously mentioned Fermi master equation. The Fermi master equation is equivalent to Fermi’s golden rule combined with the conservation of total probability [14]. This means that by introducing the concept of non-coherent evolution into quantum mechanics, we arrive at equations that already exist. Thus, we do not contradict any well-known or well-tested results of quantum mechanics. Rather, we re-derive and reinterpret them in a way that makes quantum mechanics compatible with statistical mechanics, particularly with the irreversibility found in kinetic equations. In [14], it is shown that entropy growth follows from the Fermi master equation. We will additionally show that both the equidistribution of probabilities and the Boltzmann equation can be readily derived from it.
@@ II Non-coherent transformations
We impose the non-coherence condition on the transformation, signifying that we lack knowledge of the arguments $\varphi$ of the amplitudes in Eq. (4); only their absolute values are known. To ascertain how the probabilities transform, we average over all possible values of the arguments:
$P_{k_{f}}=\frac{1}{(2\pi)^{N}}\idotsint\limits_{0}^{2\pi}\prod_{k_{i}}d\varphi_{k_{i}}\left[\sum_{k_{i}}|U_{k_{f}k_{i}}|^{2}|A_{k_{i}}|^{2}+\sum_{k^{\prime}_{i}\neq k_{i}}U^{*}_{k_{f}k^{\prime}_{i}}U_{k_{f}k_{i}}e^{i(\varphi_{k^{\prime}_{i}}-\varphi_{k_{i}})}A^{*}_{k^{\prime}_{i}}A_{k_{i}}\right].$ (5)
We see that the integral of the sum over $k^{\prime}_{i}\neq k_{i}$ equals zero since we have $\int d\varphi_{k_{i}}\exp i(\varphi_{k^{\prime}_{i}}-\varphi_{k_{i}})=0$. The integral over the first term equals $(2\pi)^{N}\sum_{\beta}|U_{\alpha\beta}|^{2}|A_{\beta}|^{2}$, since integrated function does not depend on integration variable. We see that after averaging over arguments, the column of probabilities is
$P_{k_{f}}=\sum_{\beta}|U_{k_{f}k_{i}}|^{2}|A_{k_{i}}|^{2}$
$\mathcal{P}_{f}=|\mathcal{U}|^{\circ 2}|\mathcal{A}_{i}|^{\circ 2}$
@@
Here we introduced the matrix $\mathcal{T}=|\mathcal{U}|^{\circ 2}$, the matrix of non-coherent transformation. What are the physical consequences of having the transformation law (6) instead of (3)?
@@
Although the experiment in [37] was carried out in the context of classical optics, the results can be interpreted in terms of photons and wave packets. First of all, the intensity of light is proportional to the number of photons multiplied by the energy flux associated with each, $c\hbar\omega$. When photons are reflected in a given direction, the expected number is given by the product of the number of incident photons and the probability of transmission, $NP$. By the law of large numbers, the actual number of photons does not differ significantly from the expected value. Thus, we can write $I=PNc\hbar\omega$. This allows us to use intensity and probability interchangeably.
@@
$\mathcal{P}_{f}=\int d^{N}\varphi_{n}\,...\int d^{N}\varphi_{1}\,|\mathcal{U}_{n}\Phi_{n}...\mathcal{U}_{1}\Phi_{1}\mathcal{A}_{i}|^{\circ 2}.$ (15)
@@
$P_{k_{f}}=\sum_{k_{n},k^{\prime}_{n}=1}^{N}...\sum_{k_{1},k^{\prime}_{1}=1}^{N}\sum_{k_{i},k^{\prime}_{i}=1}^{N}(U_{k_{f},k^{\prime}_{n}}...U_{k^{\prime}_{1},k^{\prime}_{i}}A_{k^{\prime}_{i}})^{*}\delta_{k^{\prime}_{n},k_{n}}...\delta_{k^{\prime}_{i},k_{i}}(U_{k_{f},k_{n}}...U_{k_{1},k_{i}}A_{k_{i}}).$ (18)
@@
$\mathcal{P}_{f}=|\mathcal{U}_{n}|^{\circ 2}...|\mathcal{U}_{1}|^{\circ 2}|\mathcal{A}_{i}|^{\circ 2}$
@@ III Non-coherent evolution
As the previous section shows, in the discrete case, the non-coherence condition, $\Delta t\gg t_{\text{coh}}$, converts a reversible unitary transformation into an irreversible stochastic transformation. In the continuous case, unitary evolution is described by the exponential of a skew-Hermitian matrix [41]. In the context of quantum theory, we typically express this as $\mathcal{U}(t)=\exp(i\mathcal{H}t)$, where $\mathcal{H}$ is a Hermitian matrix, particularly representing the Hamiltonian of the system. The continuous analog of a bistochastic transformation is a symmetric continuous-time Markov chain (CTMC). The time dependence of transition probabilities is given by $\mathcal{T}(t)=\exp(\mathcal{Q}t)$, where $\mathcal{Q}$ is the transition rate matrix (Q-matrix) [42]. Its non-diagonal elements are positive numbers representing the rates of transition between states. The diagonal elements $Q_{\alpha\alpha}$ represent the total departure rate from the state $\alpha$. They are negative, such that the sums of elements in each column are zero [^1] In probability texts, the state of the system is often represented by a row, and the transformation matrix is placed on the right of it. We use the convention adopted in physical texts: to represent the state as a column and place the transformation matrix to the left of it. This convention results in a difference in definition: in probability textbooks, the sums of elements in each row of a Q-matrix are zero.. A general Q-matrix serves as the continuous analog of a stochastic transformation, while the analog of a bistochastic transformation is a Q-matrix symmetric over the main diagonal, i.e., $Q_{\alpha\beta}=Q_{\beta\alpha}$. We are going to demonstrate that just as the unitary transformation converts into a bistochastic transformation under the condition of non-coherence, unitary evolution converts into symmetric stochastic evolution.
@@
”Let us now give a qualitative description of why this system should be irreversible. To this end, we compare it with the light beam investigated in the previous section. There, we subdivided the beam into imaginary segments of length $l_{\text{coh}}$ and found that when photons from different segments interact, the interaction is irreversible (Fig. 2). In the case of the light beam, all photons move in the same direction with the same speed, so in order for photons from different regions to interact, we had to split the beam into two parts with paths of different optical lengths.
@@
$\Psi=\sum_{k}A_{k}|k\rangle,$ (21)
where $|k\rangle$ are states of non-interacting system, and $A_{k}$ are probability amplitudes of system to be in state $|k\rangle$. Standartly we get the following equation for $A_{k}$:
@@
$\mathcal{A}(t+dt)=e^{\mathcal{M}(t)dt}\mathcal{A}(t),$
$M_{kl}(t)=-\frac{i}{\hbar}V_{kl}\exp\left(\frac{i}{\hbar}(E_{k}-E_{l})t\right),$ (24)
which is unitary transformation since $M_{kl}=M_{lk}^{\dagger}$. On a finite time interval, transformation of amplitudes can be expressed as an infinite product of exponents of the form (24), so it is unitary as a product of unitary transforms.
@@
Equation (23) is too complicated to tackle directly. To continue our investigation, we need to use an approximate method. The original Diracs approach is to employ the variation of parameters method. We choose a different approach here.
@@
$\int\limits_{0}^{\Delta t}\exp\left(\frac{i}{\hbar}(E_{k}-E_{l})t\right)A\,dt\approx A\int\limits_{0}^{\Delta t}\exp\left(\frac{i}{\hbar}(E_{k}-E_{l})t\right)dt=-\frac{i\hbar}{E_{k}-E_{l}}\left(e^{\frac{i}{\hbar}(E_{k}-E_{l})\Delta t}-1\right)A.$ (25)
We divide both sides by $\Delta t$. We define the obtained expression before $A$ in the right side of the equation as $\chi_{\Delta E}(E_{k})$, where $\Delta E=|E_{k}-E_{l}|$. We will consider two cases:
$\Delta E\Delta t\gg\hbar=>\chi_{\Delta E}(E_{k})\approx 0,$
$\Delta E\Delta t\ll\hbar=>\chi_{\Delta E}(E_{k})\approx 1.$ (26)
We see that $\chi_{\Delta E}(E_{k})$ guarantees that only states with energies close to $E_{k}$ contribute to the variation of $A_{k}$. We want to emphasize this point even more and simplify our calculation. To do that, we replace $\chi_{\Delta E}(E_{k})$ with an indicator function $\overline{\chi}_{\Delta E}(E_{k})$, which equals one in a region of energies near $E_{k}$ and zero outside of this region.
We want to find the appropriate length of this region. We will use $\chi,\overline{\chi}$ in expressions like $\int dE\chi A$, which correspond to the scalar product of amplitude, as a function of energy, over $\chi$. We want norms of $\chi,\overline{\chi}$ be equal: $\int dE|\chi|^{2}=\int dE|\overline{\chi}|^{2}$, which yields
$\Delta E\Delta t>$ $\pi\hbar=>\overline{\chi}_{\Delta E}(E_{k})=0$
$\Delta E\Delta t<$ $\pi\hbar=>\overline{\chi}_{\Delta E}(E_{k})=1.$ (27)
Later, we will see what information we have lost with this simplification.
@@
$\frac{A(\Delta t)-A(0)}{\Delta t}\approx\dot{A},$ (28)
@@
$\dot{A}_{k}=-\frac{i}{\hbar}\sum_{l}\chi_{\Delta E}(E_{k})V_{kl}A_{l}.$ (29)
Later we will not write $\chi_{\Delta E}(E_{k})$ explicitely, but will keep in mind that we sum over $l$-s such that $E_{k}-E_{l}<\Delta E$ (Fig. 4b). In the matrix form the equation is
@@
This equation is much simpler than the one we started from in (23), because the matrix elements are now time-independent. However, we note that the matrix $\mathcal{V}$ is Hermitian, so equation (30) still describes unitary evolution of the system. Thus, we have not lost this fundamental property through our simplifications.
@@
$\mathcal{A}(t)=\lim_{n\to\infty}\left(\mathcal{E}-\frac{i}{\hbar}\mathcal{V}t/n\right)^{n}\mathcal{A}(0)$ (32)
The strategy of our derivation is to take $n$ as a large number, but not instantly take the limit $n\to\infty$. If we take this limit we would get $\dot{P}=0$. Instead, we want to expand functions to the Taylor series in the precision to first non-vanishing terms in $1/n$ and analyze its asymptotic behavior at $n\to\infty$. We denote $\Delta t=t/n$. We rewrite (32) as
@@
Soon we will see the importace of keeping expression of order $\Delta t^{2}$.
We now want to analyze the non-coherent regime. This means that the state under consideration, $|k\rangle$, belongs to a sheaf of states with a width in the frequency region $\omega_{\text{idth}}$, and we choose a working timescale, $\Delta t$, such that it is much larger than the coherence time. Equivalently, we have $\Delta E\ll\omega_{\text{idth}}$ (Fig. 4b). The quantity $\Delta E$ fits into $\omega_{\text{idth}}$ many times, each time with a different phase difference for each pair of amplitudes, so we can average over the phase. This allows the formula for a series of non-coherent transformations, given in (20), to be applicable. We note that this approximation is in close analogy with the Born-Markov approximation of the Lindblad equation [16], where it is assumed that the heat bath relaxation time is small compared to the relaxation time of the system.
@@
Similarly to what we have done with unitary matrices in expression (32), we want to express bistochastic matrices $\mathcal{T}$ in a form of unity matrix plus $\mathcal{Q}\Delta t$. Using (6) we have
@@
For diagonal terms unit matrices cancel out and we have to calculate the diagonal of the matrix
$Q_{kk}=\left(\frac{\Delta t}{\hbar}\right)\left(\frac{1}{2}\mathcal{V}^{2}+\frac{1}{2}\mathcal{V}^{*^{2}}\right)_{kk}.$ (37)
For this we use the fact that $\mathcal{V}$ is Hermitian with zero diagonal elements, so we obtain
@@
We note that the first and the last expressions are equivalent, both can be used for calculation. We introduce the last form because it resembles the product rule formula. And we assume we use the last form in subsequent manipulations.
For $\mathcal{A}(\Delta t)$ we use the formula (35), setting $t=\Delta t$
$\mathcal{A}(\Delta t)=\left(\mathcal{E}-\frac{i}{\hbar}\mathcal{V}\Delta t-\frac{1}{2\hbar^{2}}\mathcal{V}^{2}\Delta t^{2}+o(\Delta t^{2})\right)\mathcal{A}(0).$ (42)
This time we want to write up the computation element-wise, to present it both ways. For $A_{k}(\Delta t)$, only keeping terms of the first order in $\Delta t$:
@@
We substitute expressions (43, 44) and their complex conjugates into (41) and imply non-coherence condition that makes $A^{*}_{k}A_{l}\rightarrow\delta_{kl}|A_{k}|^{2}=P_{k}$. Thus we obtain
@@
Taking into consideration the definition of the $\mathcal{Q}$ matrix (36, 38), we see that it is the same equation as (40). We observe that we obtain the same result with a slightly different reasoning. In the latter case, the connections with the results of the previous section are less apparent, but the differential equation arises more naturally.
@@
We begin by utilizing an approximation of a continuous spectrum. We can do it since, as we have assumed at the beginning of the section, the spectral width of the system is much greater than the energy difference between the levels of the system. Near a given state $l$, we divide the quantum numbers into two components. The first is $l_{T}$, such that small changes in its value leave the energy unchanged, so it is tangential to the energy surface. The second is $l_{N}$, where a change results in a change in energy. We express energy as a function of the quantum numbers in the form $E(l)=E(l_{0})+l_{N},\mathrm{grad}_{l}E$. Additionally, we recall expression (27) and substitute $\Delta t=\pi\hbar/\Delta E$. We then obtain:

### v6

@@ I Introduction
Another notable approach is the study of quantum entanglement and its relationship to thermodynamic irreversibility. Some researchers have proposed that the generation and disruption of entanglement within a quantum system may play a key role in the emergence of macroscopic irreversibility [31, 32, 33].
In the textbook [14], the second law of thermodynamics is derived from an equation known as the Fermi master equation. However, the equation itself is noted to lack a proper derivation. A modern justification for this idea is provided by the Lindblad master equation [15], which describes systems weakly coupled to their environments. In particular, the Fermi master equation is regarded as a Born-Markov approximation of the Lindblad equation [16]. These equations have been shown to exhibit irreversible behavior, even when the underlying microscopic dynamics are unitary [17]. This approach has been applied to various models in quantum thermodynamics [18, 19, 20, 21]. A particularly important result of this line of research is the identification of conditions under which systems reach a steady state resembling the single-particle distribution functions (Maxwell-Boltzmann, Fermi-Dirac or Bose-Einstein distributions) [22, 23, 24].
Analogously, in Ref. [34], the authors derive the second law of thermodynamics from a quantum-mechanical analysis of a total system partitioned into an object and its environment. This type of theory has attracted considerable attention in recent years [35, 36, 37]. Additionally, the theory of non-Markovian quantum evolution [38, 39, 40, 41] has garnered significant interest. It focuses on cases where the system’s dynamics are influenced by interactions with the external world.
@@
After performing the averaging over arguments, we obtain the equation of non-coherent evolution. We find that it coincides with the previously mentioned Fermi master equation. The Fermi master equation is equivalent to Fermi’s golden rule combined with the conservation of total probability [14]. This means that by introducing the concept of non-coherent evolution into quantum mechanics, we arrive at equations that already exist. Thus, we do not contradict any well-known or well-tested results of quantum mechanics. Rather, we re-derive and reinterpret them in a way that makes quantum mechanics compatible with statistical mechanics, particularly with the irreversibility found in kinetic equations. In [14], it is shown that entropy growth follows from the Fermi master equation. We will additionally show that both the equidistribution of probabilities and the Boltzmann equation in important particular cases can be readily derived from it.
@@ II Non-coherent transformations
We impose the non-coherence condition on the transformation, signifying that we lack knowledge of the arguments $\varphi$ of the amplitudes in Eq. (4); only their absolute values are known. To ascertain how the probabilities transform, we average over all possible values of the arguments. To distinguish the probability calculated with non-coherence conditions imposed from the one calculated ”normally” by formula (4), we use a tilde:
$\widetilde{P_{k_{f}}}=\frac{1}{(2\pi)^{N}}\idotsint\limits_{0}^{2\pi}\prod_{k_{i}}d\varphi_{k_{i}}\left[\sum_{k_{i}}|U_{k_{f}k_{i}}|^{2}|A_{k_{i}}|^{2}+\sum_{k^{\prime}_{i}\neq k_{i}}U^{*}_{k_{f}k^{\prime}_{i}}U_{k_{f}k_{i}}e^{i(\varphi_{k^{\prime}_{i}}-\varphi_{k_{i}})}A^{*}_{k^{\prime}_{i}}A_{k_{i}}\right].$ (5)
We see that the integral of the sum over $k^{\prime}_{i}\neq k_{i}$ equals zero since we have $\int d\varphi_{k_{i}}\exp i(\varphi_{k^{\prime}_{i}}-\varphi_{k_{i}})=0$. The integral over the first term equals $(2\pi)^{N}\sum_{k_{i}}|U_{k_{f}k_{i}}|^{2}|A_{k_{i}}|^{2}$, since integrated function does not depend on integration variable. We see that after averaging over arguments, the column of probabilities is
$\widetilde{P_{k_{f}}}=\sum_{k_{i}}|U_{k_{f}k_{i}}|^{2}|A_{k_{i}}|^{2}$
$\widetilde{\mathcal{P}_{f}}=|\mathcal{U}|^{\circ 2}|\mathcal{A}_{i}|^{\circ 2}.$ (6)
When there are no complex-valued variables on the right-hand side of the equation, we can omit the tilde on the left, since in this case we assume non-coherence without any possible confusion. We introduce the matrix $\mathcal{T}=|\mathcal{U}|^{\circ 2}$, the matrix of non-coherent transformation and rewrite the formula as:
@@
What are the physical consequences of having the transformation law (7) instead of (3)?
@@
Although the experiment in [43] was carried out in the context of classical optics, the results can be interpreted in terms of photons and wave packets. First of all, the intensity of light $I$ is proportional to the number of photons $N$ multiplied by the energy flux associated with each. The flux associated with each photon is its velocity, which is the speed of light, multiplied by its energy, given by Planck’s formula: $c\hbar\omega$. When photons are reflected in a given direction, the expected number is given by the product of the number of incident photons and the probability of transmission, $NP$. By the law of large numbers, the actual number of photons does not differ significantly from the expected value. Thus, we can write $I=PNc\hbar\omega$. This allows us to use intensity and probability interchangeably. We note that $P$ and $N$ are dimensionless quantities, while intensity and energy flux are synonymous and have the same dimension, so the dimensions of the left-hand side and the right-hand side of the formula are the same.
@@
$\widetilde{\mathcal{P}_{f}}=\int d^{N}\varphi_{n}\,...\int d^{N}\varphi_{1}\,|\mathcal{U}_{n}\Phi_{n}...\mathcal{U}_{1}\Phi_{1}\mathcal{A}_{i}|^{\circ 2}.$ (16)
@@
$\widetilde{P_{k_{f}}}=\sum_{k_{n},k^{\prime}_{n}=1}^{N}...\sum_{k_{1},k^{\prime}_{1}=1}^{N}\sum_{k_{i},k^{\prime}_{i}=1}^{N}(U_{k_{f},k^{\prime}_{n}}...U_{k^{\prime}_{1},k^{\prime}_{i}}A_{k^{\prime}_{i}})^{*}\delta_{k^{\prime}_{n},k_{n}}...\delta_{k^{\prime}_{i},k_{i}}(U_{k_{f},k_{n}}...U_{k_{1},k_{i}}A_{k_{i}}).$ (19)
@@
$\widetilde{\mathcal{P}_{f}}=|\mathcal{U}_{n}|^{\circ 2}...|\mathcal{U}_{1}|^{\circ 2}|\mathcal{A}_{i}|^{\circ 2}$
@@ III Non-coherent evolution
As the previous section shows, in the discrete case, the non-coherence condition, $\Delta t\gg t_{\text{coh}}$, converts a reversible unitary transformation into an irreversible stochastic transformation. In the continuous case, unitary evolution is described by the exponential of a skew-Hermitian matrix [47]. In the context of quantum theory, we typically express this as $\mathcal{U}(t)=\exp(i\mathcal{H}t)$, where $\mathcal{H}$ is a time-independent Hermitian matrix, particularly representing the Hamiltonian of the system. The continuous analog of a bistochastic transformation is a symmetric continuous-time Markov chain (CTMC). The time dependence of transition probabilities is given by $\mathcal{T}(t)=\exp(\mathcal{Q}t)$, where $\mathcal{Q}$ is the transition rate matrix (Q-matrix) [48]. Its non-diagonal elements are positive numbers representing the rates of transition between states. The diagonal elements $Q_{\alpha\alpha}$ represent the total departure rate from the state $\alpha$. They are negative, such that the sums of elements in each column are zero [^1] In probability texts, the state of the system is often represented by a row, and the transformation matrix is placed on the right of it. We use the convention adopted in physical texts: to represent the state as a column and place the transformation matrix to the left of it. This convention results in a difference in definition: in probability textbooks, the sums of elements in each row of a Q-matrix are zero.. A general Q-matrix serves as the continuous analog of a stochastic transformation, while the analog of a bistochastic transformation is a Q-matrix symmetric over the main diagonal, i.e., $Q_{\alpha\beta}=Q_{\beta\alpha}$. We are going to demonstrate that just as the unitary transformation converts into a bistochastic transformation under the condition of non-coherence, unitary evolution converts into symmetric stochastic evolution.
@@
Let us now give a qualitative description of why this system should be irreversible. To this end, we compare it with the light beam investigated in the previous section. There, we subdivided the beam into imaginary segments of length $l_{\text{coh}}$ and found that when photons from different segments are incident on the interface simultaneously, the transformation of associated energy flux is irreversible (Fig. 2). In the case of the light beam, all photons move in the same direction with the same speed, so in order for photons from different regions to meet, we had to split the beam into two parts with paths of different optical lengths.
@@
$|\Psi\rangle=\sum_{k}A_{k}|k\rangle,$ (22)
where $|k\rangle$ are states of non-interacting system, and $A_{k}$ are probability amplitudes of system to be in state $|k\rangle$. We get the following equation for $A_{k}$:
@@
$\mathcal{A}(t+dt)=e^{\mathcal{M}(t)dt}\mathcal{A}(t),$ (25)
where elements of matrix $\mathcal{M}(t)$ are defined by expression in the right side of (24):
$M_{kl}(t)=-\frac{i}{\hbar}V_{kl}\exp\left(\frac{i}{\hbar}(E_{k}-E_{l})t\right).$ (26)
We see that evolution described by Eq. (25) is unitary since $M_{kl}$ is skew-Hermitian: $M_{kl}=-M_{lk}^{\dagger}$. On a finite time interval, transformation of amplitudes can be expressed as an infinite product of exponents of the form (25), so it is unitary as a product of unitary transforms.
@@
Equation (24) is too complicated to tackle directly. To continue our investigation, we need to use an approximate method. The original Dirac’s approach is to employ the variation of parameters method. We choose a different approach here.
@@
$\int\limits_{0}^{\Delta t}\exp\left(\frac{i}{\hbar}(E_{k}-E_{l})t\right)A_{l}\,dt\approx A_{l}\int\limits_{0}^{\Delta t}\exp\left(\frac{i}{\hbar}(E_{k}-E_{l})t\right)dt=-\frac{i\hbar}{E_{k}-E_{l}}\left(e^{\frac{i}{\hbar}(E_{k}-E_{l})\Delta t}-1\right)A_{l}.$ (27)
We divide both sides by $\Delta t$. We want to take a closer look at the obtained dimensionless expression before $A$ on the right-hand side of the equation. We define
$\chi_{kl}=-\frac{i\hbar}{(E_{k}-E_{l})\Delta t}\left(e^{\frac{i}{\hbar}(E_{k}-E_{l})\Delta t}-1\right).$ (28)
This expression has two limiting cases:
$|E_{k}-E_{l}|\gg\hbar/\Delta t=>\chi_{kl}(E_{k})\approx 0,$
$|E_{k}-E_{l}|\ll\hbar/\Delta t=>\chi_{kl}(E_{k})\approx 1.$ (29)
We see that $\chi_{kl}$ guarantees that only states with energies close to $E_{k}$ contribute to the variation of $A_{k}$. Physically, it is connected with the time–energy uncertainty principle: we cannot distinguish energies that differ by less than $\Delta E$ over a time shorter than $\hbar/\Delta E$.
We want to emphasize this point even more and simplify our calculation. To do that, we replace $\chi_{kl}$ with an indicator function $\overline{\chi}_{kl}$, which equals one in a region of energies near $E_{k}$ and zero outside of this region. We call the energy radius $\Delta E$:
$|E_{k}-E_{l}|>\Delta E$ $=>\overline{\chi}_{kl}=0$
$|E_{k}-E_{l}|<\Delta E$ $=>\overline{\chi}_{kl}=1.$ (30)
We want to choose the value of $\Delta E$ so that
$\int dE|\chi_{kl}|^{2}=\int dE|\overline{\chi}_{kl}|^{2}.$ (31)
Introducing integration variable $E=(E_{k}-E_{l})$
$\int dE|\chi_{kl}|^{2}=\frac{4\hbar^{2}}{\Delta t^{2}}\int dE\frac{\sin^{2}(E\Delta t/2\hbar)}{E^{2}}=2\pi\hbar/\Delta t.$ (32)
From the definition of the indicator function
$\int dE|\overline{\chi}_{kl}|^{2}=\int dE|\overline{\chi}|=\int_{-\Delta E}^{\Delta E}dE=2\Delta E.$ (33)
Substituting Eqs. (32, 33) to (31) we find that $\Delta E=\pi\hbar/\Delta t$. Comparing this with Eq. (30) we get
$|E_{k}-E_{l}|>\pi\hbar/\Delta t$ $=>\overline{\chi}_{kl}=0$
$|E_{k}-E_{l}|<\pi\hbar/\Delta t$ $=>\overline{\chi}_{kl}=1.$ (34)
Condition (31) is hard to justify at this point. At the end of the derivation, we will see why it is necessary; see the discussion after Eq. (56). There, we explain why the exchange of $\chi_{kl}$ with $\overline{\chi}_{kl}$ is possible and what information we have lost in doing so.
@@
$\frac{A_{k}(\Delta t)-A_{k}(0)}{\Delta t}\approx\dot{A}_{k},$ (35)
@@
$\dot{A}_{k}=-\frac{i}{\hbar}\sum_{l}\overline{\chi}_{kl}V_{kl}A_{l}.$ (36)
Later we will not write $\overline{\chi}_{kl}$ explicitely, but will keep in mind that we sum over $l$-s such that $E_{k}-E_{l}<\Delta E$ (Fig. 4b). In the matrix form the equation is
@@
This equation is much simpler than the one we started from in (24), because the matrix elements are now time-independent. However, we note that the matrix $\mathcal{V}$ is Hermitian, as we can see from the explicit form (36), so equation (37) still describes unitary evolution of the system. Thus, we have not lost this fundamental property through our simplifications.
@@
$\mathcal{A}(t)=\lim_{n\to\infty}\left(\mathcal{E}-\frac{i}{\hbar}\mathcal{V}t/n\right)^{n}\mathcal{A}(0).$ (39)
Here $\mathcal{E}$ denotes the identity matrix.
The strategy of our derivation is to take $n$ as a large number, but not instantly take the limit $n\to\infty$. If we take this limit we would get $\dot{P}=0$. Instead, we want to expand the exponent in Eq. (38) in a Taylor series, keeping terms up to the first non-vanishing order in $1/n$, so we obtain a small change over the time $t/n$:
$\mathcal{A}(\Delta t)=\left(\mathcal{E}-\frac{i}{\hbar}\mathcal{V}\Delta t-\frac{1}{2\hbar^{2}}\mathcal{V}^{2}\Delta t^{2}+O(\Delta t^{3})\right)\mathcal{A}(0),$ (40)
Where we denote $\Delta t=t/n$. Soon we will see the importace of keeping expression of order $\Delta t^{2}$.
Then we multiply $n$ small changes to obtain the evolution over an arbitrary time interval $t$:
@@
In the end we will analyze the asymptotic behavior of obtained expression at $n\to\infty$. In this way, we have rewritten formula (38).
We now want to analyze the non-coherent regime. This means that the state under consideration, $|k\rangle$, belongs to a sheaf of states with a width in the frequency region $\omega_{\text{idth}}$, and we choose a working timescale, $\Delta t$, such that it is much larger than the coherence time. Equivalently, we have $\Delta E\ll\hbar\omega_{\text{idth}}$ (Fig. 4b). The quantity $\Delta E$ fits into $\hbar\omega_{\text{idth}}$ many times, each time with a different phase difference for each pair of amplitudes, so we can average over the phase. This allows the formula for a series of non-coherent transformations, given in (21), to be applicable. We note that this approximation is in close analogy with the Born-Markov approximation of the Lindblad equation [16], where it is assumed that the heat bath relaxation time is small compared to the relaxation time of the system.
@@
Similarly to what we have done with unitary matrices in expression (39), we want to express bistochastic matrices $\mathcal{T}$ in a form of unit matrix plus $\mathcal{Q}\Delta t$. Using (7) we have
@@
For diagonal terms the identity matrices in Eq. (43) cancel out, and we have to calculate the diagonal of the matrix
$Q_{kk}=\left(\frac{\Delta t}{\hbar^{2}}\right)\left(\frac{1}{2}\mathcal{V}^{2}+\frac{1}{2}\mathcal{V}^{*^{2}}\right)_{kk},$ (45)
which is obtained by multiplying the second-order terms in $\Delta t$ from Eq. (40) with the identity matrices. For this we use the fact that $\mathcal{V}$ is Hermitian with zero diagonal elements, so we obtain
@@
We note that the first and the last expressions are equivalent, both can be used for calculation. We introduce the last form because it resembles the product rule formula. And we are going to use the last form in subsequent manipulations.
For $\mathcal{A}(\Delta t)$ we use the formula (40). This time we want to write up the computation element-wise, to present it both ways. For $A_{k}(\Delta t)$, only keeping terms of the first order in $\Delta t$:
@@
We substitute expressions (50, 51) and their complex conjugates into (49) and imply non-coherence condition that makes $A^{*}_{k}A_{l}\rightarrow\delta_{kl}|A_{k}|^{2}=P_{k}$. Thus we obtain
@@
Taking into consideration the definition of the $\mathcal{Q}$ matrix (44, 46), we see that it is the same equation as (48). We observe that we obtain the same result with a slightly different reasoning. In the latter case, the connections with the results of the previous section are less apparent, but the differential equation arises more naturally.
@@
We begin by utilizing an approximation of a continuous spectrum. We can do it since, as we have assumed at the beginning of the section, the spectral width of the system is much greater than the energy difference between the levels of the system. Expression $E(l)=\text{const}$, defines an energy surface in the space of quantum numbers $l$. Near a given state $l$, we divide the quantum numbers into two components. The first component, $l_{T}$, is tangential to the energy surface; in other words, small changes in its value leave the energy unchanged. The second component, $l_{N}$, is normal to the energy surface. Additionally, we recall expression (34) and substitute $\Delta t=\pi\hbar/\Delta E$ into the second term on the right-hand side of Eq. (52). We obtain: